pls help me ,I couldnt solve these


q1: if y=ex+e^x+e^x+...........∞,then dy/dx equals


q2:if x=at2 and y=2at then (xy*d2y/dx2)  is equal to


 


sir as u asked i checked both the ques , both are from the course material and are correct


 the ans for q1 is = y/(1-y) .


 sir in q2 could u pls exp how the chain rule is applied in diff. second time


thank u

Grade:10

1 Answers

piyush shukla
13 years ago

let y=e^x+e^x+...........∞,

then y=ex+y

now diff

  dy/dx=ex+y(1+ dy/dx)

dy/dx(1- ex+y)= ex+y

dy/dx=( ex+y)/( 1- ex+y)

dy/dx=y/(1-y)